# Mitigating Attention Hacking in Preference-Based Reward Modeling via Interaction Distillation

### v1

@@ Introduction
OOD Preference Perception.
Firstly, we train reward models on the HH-RLHF dataset and evaluate them on both OOD reward model benchmarks. Table 2 reports that Id-Rm achieves superior performance compared to other reward modeling methods on RMB’s two tasks and Reward Bench’s Chat-Hard and Reasoning tasks. With average accuracy of 69.92%, the results demonstrate that Id-Rm, by modeling more comprehensive attention mechanisms, not only adjusts attention allocation within sequences but also sensitively captures token-level interactions between preference pairs. This enables the learning of more generalizable information and improves accuracy in recognizing out-of-distribution preference pairs.
Ablation & Compatibility Study
@@
Compatibility Study.
We employ three easily integrable algorithms: adaptive margining, label flipping, and weight averaging (Burns et al. 2024; Touvron et al. 2023; Wang et al. 2024b; Rame et al. 2024) for compatibility experiments. As shown in Table 3, the RM’s performance achieves further improvement when integrating these algorithms, with weight averaging yielding the most significant boost. This demonstrates that interactive distillation effectively maintains compatibility with other optimization algorithms without altering the original model’s architecture.
Hyperparameter Sensitivity Study
@@
Related Work
Reward Modeling in RLHF
@@
Reinforcement learning from human feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022) provides a crucial and straightforward approach to align large language models (LLMs) with human preferences, constituting an essential training procedure for current mainstream LLMs (Achiam et al. 2023; Touvron et al. 2023; Yang et al. 2024; Guo et al. 2025) and improving model utility, honesty, and harmlessness (Ethayarajh, Choi, and Swayamdipta 2022).
@@
As the core component of RLHF, the reward model (RM) provides critical feedback for responses sampled by the policy model (Dou et al. 2024; Wang et al. 2024a; Dou et al. 2025b). However, reward hacking (Miao et al. 2024; Fu et al. 2025) poses a critical challenge for current reward models, as reward models are often hacked by biases in the data, leading to erroneous judgments. While existing optimization approaches primarily focus on addressing noise in the data (Touvron et al. 2023; Burns et al. 2024; Rame et al. 2024; Miao et al. 2024), they largely overlook the inadequate attention mechanisms in the preference modeling framework. We pioneer the investigation of attention hacking stemming from structural defects in reward models.
Knowledge Distillation for Language Models
@@
The knowledge distillation techniques for large language models have evolved from simple output imitation to a series of sophisticated methodologies. Among them, a foundational and widely applied technique is supervised fine-tuning (SFT), which involves training the student model using prompt-response data pairs generated by the teacher model (Taori et al. 2023). The success of such methods largely depends on the quality and diversity of the training data. To this end, the research community has explored various data generation strategies: extensions of self-instruct (Wang et al. 2022), which bootstrap generation from a small seed set of instructions, while more advanced approaches like Evol-Instruct evolve initial instructions by increasing their complexity and depth to enhance the model’s reasoning capabilities (Xu et al. 2023).
@@
For white-box models where internal parameters are accessible, feature-based distillation methods aim to minimize the differences between the teacher and student models in terms of output probability distributions (Gu et al. 2024b). Furthermore, a cutting-edge direction is self-improvement, where models iteratively refine and enhance their knowledge through frameworks like self-play (Chen et al. 2024) or self-rewarding mechanisms (Yuan et al. 2024). However, this study is the first to propose interaction distillation specifically for reward models, achieving knowledge distillation at the attention level.
Attention Patterns in Language Models
The attention mechanism in neural networks has been a widely discussed topic (Zang and Liu 2023b, 2024b; Mo et al. 2025). Research indicates that pre-trained language models (PLMs) based on Transformer architectures exhibit relatively static attention patterns (Gu et al. 2024a). Specifically, these models tend to focus on fixed text segments regardless of input variations, potentially overlooking crucial information at other positions (Dou et al. 2025a). These phenomena have been extensively observed across various applications including comprehension tasks (Zang and Liu 2023a; Stacey, Belinkov, and Rei 2022), long-text generation (Ye et al. 2025), and model evaluation (Zang et al. 2025). For the first time, we analyze the attention hacking phenomenon in preference-based reward modeling, which stems from: (1) the decoder-only architectures, and (2) the independent Siamese-encoding paradigm.
Conclusion
@@
We propose a novel solution termed “Interaction Distillation” to address the “attention hacking” in preference-based reward models. This approach introduces an interaction-based NLU model to distill its comprehensive attention patterns into reward models. Remarkably, it effectively rectifies the model’s inadequate token-level interaction via an attentional distillation objective, without introducing inference costs. Compared to reward modeling optimization methods that primarily address data noise, our approach generates more stable and generalizable reward signals. This suggests we can move beyond merely mitigating reward model bias at the data noise level, instead capture the intrinsic structure of complex human preferences through attention mechanisms.

### v3

@@ 4.2.1 RLHF tasks
4.2.2 OOD Preference Perception
Firstly, we train reward models on the HH-RLHF dataset and evaluate them on both OOD reward model datasets and more comprehensive reward modeling benchmarks. Table 2 reports that Id-Rm achieves superior performance compared to other reward modeling methods on RMB’s two tasks and Reward Bench’s Chat-Hard and Reasoning tasks. With average accuracy of 68.25%, the results demonstrate that Id-Rm, by modeling more comprehensive attention mechanisms, not only adjusts attention allocation within sequences but also sensitively captures token-level interactions between preference pairs. This enables the learning of more generalizable information and improves accuracy in recognizing out-of-distribution preference pairs.
4.3 Ablation & Compatibility Study
@@
4.3.2 Compatibility Study
We employ three easily integrable algorithms: adaptive margining, label flipping, and weight averaging Burns et al. (2024); Touvron et al. (2023); Wang et al. (2024b); Rame et al. (2024) for compatibility experiments. As shown in Table 3, the RM’s performance achieves further improvement when integrating these algorithms, with weight averaging yielding the most significant boost. This demonstrates that interactive distillation effectively maintains compatibility with other optimization algorithms without altering the original model’s architecture.
4.4 Hyperparameter Sensitivity Study
@@ 4.4.2 Hyperparameter $K$
5 Related Work
5.1 Reward Modeling in RLHF
@@
Reinforcement learning from human feedback  Ouyang et al. (2022); Bai et al. (2022) provides a crucial and straightforward approach to align large language models with human preferences, constituting an essential training procedure for current mainstream LLMs Achiam et al. (2023); Touvron et al. (2023); Yang et al. (2024); Guo et al. (2025) and improving model utility, honesty, and harmlessness Ethayarajh et al. (2022).
@@
As the core component of RLHF, the reward model provides critical feedback for responses sampled by the policy model Dou et al. (2024); Wang et al. (2024a); Dou et al. (2025). The paradigm of reward modeling can be divided into generative reward models Mahan et al. (2024); Zhang et al. (2024); Liu et al. (2025) and discriminative reward models. However, discriminative reward hacking Miao et al. (2024); Fu et al. (2025) poses a critical challenge for current reward models, as reward models are often hacked by biases in the data, leading to erroneous judgments. While existing optimization approaches primarily focus on addressing noise in the data Touvron et al. (2023); Burns et al. (2024); Rame et al. (2024); Miao et al. (2024), they largely overlook the inadequate attention mechanisms in the preference modeling framework. We pioneer the investigation of attention hacking stemming from structural defects in reward models.
5.2 Knowledge Distillation for Language Models
@@
The knowledge distillation techniques for large language models have evolved from simple output imitation to a series of sophisticated methodologies. Among them, a foundational and widely applied technique is supervised fine-tuning (SFT), which involves training the student model using prompt-response data pairs generated by the teacher model Taori et al. (2023). The success of such methods largely depends on the quality and diversity of the training data. To this end, the research community has explored various data generation strategies: extensions of self-instruct Wang et al. (2022), which bootstrap generation from a small seed set of instructions, while more advanced approaches like Evol-Instruct evolve initial instructions by increasing their complexity and depth to enhance the model’s reasoning capabilities Xu et al. (2023).
@@
For white-box models where internal parameters are accessible, feature-based distillation methods aim to minimize the differences between the teacher and student models in terms of output probability distributions Gu et al. (2024). Furthermore, a cutting-edge direction is self-improvement, where models iteratively refine and enhance their knowledge through frameworks like self-play Chen et al. (2024) or self-rewarding mechanisms Yuan et al. (2024). However, this study is the first to propose interaction distillation specifically for LLM based reward models, achieving knowledge distillation at the attention level.
6 Conclusion
@@
We propose a novel solution termed “Interaction Distillation” to address the “attention hacking” in discriminative reward models. This approach introduces an interaction-based NLU model to distill its comprehensive attention patterns into reward models. Remarkably, it effectively rectifies the model’s inadequate token-level interaction via an attentional distillation objective, without introducing inference costs. Compared to reward modeling optimization methods that primarily address data noise, our approach generates more stable and generalizable reward signals. This suggests we can move beyond merely mitigating reward model bias at the data noise level, instead capture the intrinsic structure of complex human preferences through attention mechanisms.